OpenStudy (anonymous):

The first four terms of a sequence are shown below. 9, 5, 1, -3 Which of the following functions best defines this sequence? f(1) = 9, f(n + 1) = f(n) - 4; for n ≥ 1 f(1) = 9, f(n + 1) = f(n) + 4; for n ≥ 1 f(1) = 9, f(n + 1) = f(n) - 5; for n ≥ 1 f(1) = 9, f(n + 1) = f(n) + 5; for n ≥ 1

OpenStudy (anonymous):

How the heck am I supposed to solve a function where, for example, f(n)-4=f(n+1) What does the = signify? We never learned this. I can solve regular recursive functions.

OpenStudy (anonymous):

f(1) = 9, f(n + 1) = f(n) - 4; for n ≥ 1 would be your answer Dave. f(n)-4=f(n+1) because it is the nth term. it is basically subtracting 4 from ever term your on and equals the next.

OpenStudy (phi):

f(n+1)= f(n) -4 is a "rule" say n is 1, then this says f(1+1) = f(1) - 4 f(1) is the first term in the sequence, which you see is 9 f(2) = 9-4 f(2)= 5 did the rule work? If so, that is the answer

OpenStudy (anonymous):

after you subtract four you just subtract it again to get your next term

OpenStudy (anonymous):

Ahhh, i get it now.

OpenStudy (phi):

the n stands for which number in the sequence... starting with 1

OpenStudy (anonymous):

thanks all can turn back time but i cant solve basic recursive functions

OpenStudy (phi):

f(n) stands for the actual number, n for the "place" so n=3 means the 3rd number in the list f(3) means the value of that 3rd number.

OpenStudy (anonymous):

lol at least math is better than being turned into a frog